OpenStudy (anonymous):

Suppose that a guitar string has a length of 1.1 m, a mass of 0.12 kg and a tension of 125 N. What is the speed of the wave on the string?

OpenStudy (apoorvk):

wave speed 'v' ---> \[v = \frac{\sqrt{T}} {\sqrt{mu}}\] where, T=tension mu=linear mass density (i.e. mass per unit length of the wire)

OpenStudy (apoorvk):

that means, mu=mass/length

OpenStudy (anonymous):

Hi you can use from this: \[v=\sqrt{T/\lambda }\] where lambda is mass per meter